OpenStudy (anonymous):

Use properties of real numbers to rewrite the expresion. (3/5) (5/3)

OpenStudy (zzr0ck3r):

(3/5)(5/3)= (3*5)/(5*3) = (3*5)/(3*5) = (3/3)*(5/5) = 1*1 = 1